# and Clade-Specific Deviations

### 27 Mar 2026 v1

@@ 1  Introduction
1.1  Time is relative: the shrew and the elephant
Time is not experienced equally by all living organisms. A pygmy shrew (Suncus etruscus, $\approx 2\,\text{g}$) burns through life at 835 heartbeats per minute and dies within two years; an African elephant ($\approx 4{,}000\,\text{kg}$) beats its heart at 28 beats per minute and endures it for seven decades. To an external observer, their chronological lifetimes differ by a factor of thirty-five. Yet when duration is measured not by the calendar but by the organism’s own internal rhythm—the cumulative count of heartbeats—the two animals are, in a deep sense, contemporaries. Each accumulates close to $10^{9}$ cardiac cycles before death; that is,
@@ 1.1  Time is relative: the shrew and the elephant
where $f_{H}$ is the resting heart rate and $L$ is the maximum lifespan.
This near-equality is not a coincidence. It reflects a thermodynamic truth about living systems: an organism does not age by years but by physiological cycles, and the total number of such cycles it may sustain is bounded by a finite dissipative budget. Precisely as Einstein’s special relativity teaches that proper time is intrinsic to each observer and cannot be universalised, biological proper time is intrinsic to each organism and cannot be read from a wall clock [12]. A hummingbird’s minute is dense with entropy production; a tortoise’s minute is sparse; yet both traverse the same arc of internal duration toward the same thermodynamic terminus.
The dimensionless product
$\ell_{i}\;\equiv\;\log_{10}\!\bigl(f_{H,i}\cdot L_{i}\cdot 525{,}960\bigr)\;\approx\;9.06$ (2)
@@
captures this empirical regularity, where $f_{H,i}$ (beats min-1) is the resting heart rate, $L_{i}$ (yr) is the maximum natural lifespan, and $525{,}960\,\text{min\,yr}^{-1}$ is the conversion factor. Equivalently, $N_{\star}=f_{H,i}\,L_{i}\cdot 525{,}960\approx 10^{9}$ cardiac cycles per lifetime. First noted through the constancy of mass-specific lifetime energy expenditure by Rubner [1], and quantified as a universal temporal allometry by Lindstedt and Calder [2]—with the heartbeat count itself explicitly computed by Livingstone and Kuehn [37] and later by Levine [38]—this invariant has since accumulated overwhelming empirical support across body masses spanning ten orders of magnitude—from the two-gram Etruscan shrew to a four-tonne elephant.
1.2  What prior theories have left unresolved
Despite its empirical solidity, $N_{\star}\approx 10^{9}$ has never received a thermodynamic motivation from physical principles, nor a predictive framework for the systematic departures from it observed in the specific clades.
Allometric accounts.
The West–Brown–Enquist (WBE) fractal vascular network theory [3] derives $f_{H}\propto M^{-1/4}$ and $L\propto M^{+1/4}$ from network optimisation, predicting $f_{H}L\propto M^{0}$ by exponent cancellation, where $M$ is the body mass. This kinematic argument elegantly explains why the product is mass-independent within any clade sharing the WBE geometry, but it does not predict the numerical value $10^{9}$, nor the systematic departures of primates, bats, and birds from the mammalian baseline. All three clades share the same vascular allometry yet differ from each other and from the baseline by factors of two to eight in $N_{\star}$. We show below (Section 9.3) that the WBE kinematic null is rejected at the clade level ($F=12.7$, $p<0.001$).
Energetic accounts.
[…]
Metabolic-level boundaries.
Glazier [5] has shown that metabolic scaling exponents vary systematically with metabolic level, taxon, and life-history strategy. The PBTE framework accommodates this: the invariant depends not on any individual exponent but on two exponents summing to zero—a condition preserved even when each deviates from the canonical $\pm 1/4$, provided their sum remains near zero. Glazier’s metabolic-level hypothesis thus motivates, rather than undermines, the clade-multiplier formalism of Section 5.
1.3  What is new in this work
The present paper makes four contributions that, to our knowledge, have not appeared previously in the literature.
  1. 1.
    We derive $N_{\star}\approx 10^{9}$ from the non-equilibrium second law by treating the adult organism as a metabolic non-equilibrium steady state (NESS) and introducing the empirical closure $\dot{e}_{p,i}=\sigma_{0,i}f_{i}$.
  2. 2.
    We show that the correct primitive invariant is $N_{\star}$, not lifetime energy per unit mass; the latter is a derived consequence valid only when body temperature and the mass-specific entropy cost per cycle are both approximately constant.
  3. 3.
    We place the invariant in a fully falsifiable framework: explicit assumptions with stated empirical status, a domain-of-validity table, and five numerical falsification criteria (Section 12).
  4. 4.
    We derive, from a unified thermodynamic multiplier $\Phi_{C}=\Phi_{\rm duty}\cdot\Phi_{\rm thermal}\cdot\Phi_{\rm mito+oxid}\cdot\Phi_{\rm haz}$, the quantitative longevity deviations of four clades—primates, bats, birds, and cetaceans—from independently measured physiology. Each factor is calibrated to clade-specific physiological measurements; no factor is fitted directly to lifespan data.
Two contributions deserve particular emphasis in this study. For primates, the elevated lifetime count $\langle N_{\star}\rangle_{\rm prim}\approx(2\text{--}3)\times 10^{9}$ is derived from a neuro-metabolic entropy model: investment of a larger fraction of resting metabolism in neural tissue reduces the entropy produced per cardiac cycle, expanding the effective cycle budget. Prior treatments noted this deviation descriptively but offered no causal mechanisms. For bats, the extreme longevity ($\Phi_{\rm bat}\approx 7.9\times N_{0}$) arises from the product of a duty-cycle factor (cardiac suppression during torpor) and an Arrhenius thermal factor during hibernation—two mechanisms that act simultaneously and multiplicatively, a synergy not previously given a thermodynamic motivation.
The thermodynamic framework, the unified multiplier formalism, and the clade corrections for primates, bats, birds, and cetaceans are all new to this work.
1.4  Structure of this paper
@@
Section 2 establishes the notation. Section 3 derives the PBTE invariant from first principles. Section 4 defines biological proper time and its relationship to ageing. Section 5 derives the generalised clade-multiplier framework. Section 6 applies the framework to primates, bats, birds, and cetaceans, with fully worked numerical examples. Section 8 describes the comparative dataset of 230 species. Section 9 presents the statistical tests used in this study. Section 10 delimits the domain of the applicability. Section 11 discusses implications and the outstanding calorimetric experiment that would convert PBTE from a statistically supported regularity to a fully tested conservation law. Section 12 states explicit falsification criteria.
@@ 8.3  Arrhenius correction for ectotherms
with $E_{a}=0.65$ eV and $T_{\rm ref}=310$ K [11]. This correction removes approximately 75% of the raw gap between reptiles and the mammalian baseline.
@@ 9  Results
Figure 1: Lifetime cycle count $\ell$ across endothermic vertebrates. (a) Log-log plot of maximum lifespan $L$ (years) vs resting heart rate $f_{H}$ (bpm) for 230 vertebrate species. Non-primate placentals (filled circles), primates (triangles), birds (squares), bats (diamonds), cetaceans (open circles), and Arrhenius-corrected ectotherms (crosses). OLS line on non-primate placentals (slope $0.90\pm 0.06$, black solid); PBTE null slope 1-1 (dashed). Diagonal iso-$\ell$ contours at $\ell=8,9,10$. (b) Notched box-plots of $\ell$ by clade. Horizontal dashed line at non-primate mammalian mean $\bar{\ell}=8.995$. [Figure generated from species data in Extended Data Tables 1–8.]
@@ 9.1  Cross-clade regression
Figure 1 displays the full 230-species dataset. Panel (a) shows the log-log scatter of maximum lifespan $L$ against resting heart rate $f_{H}$; the OLS line is fitted to the 43 directly-measured non-primate placentals and iso-$\ell$ contours are drawn at $\ell=8$, 9, 10. Panel (b) shows clade distributions of $\ell$ as notched box-plots; the mammalian baseline $\bar{\ell}=8.995$ is the dashed line.
OLS on the $n=43$ non-primate placentals with directly measured (non-imputed) heart rates: $\hat{\beta}=-0.90\pm 0.06$ (s.e.), $R^{2}=0.86$, $F$-test $p=0.09$ against $\beta=-1$. The intercept gives $\bar{\ell}=8.995\pm 0.025$, corresponding to $N_{0}\approx 1.02\times 10^{9}$. Repeating the OLS on all $n=46$ non-primate placentals (including 3 species with allometrically imputed heart rates) yields $\hat{\beta}=-0.91\pm 0.06$, $p=0.13$ against $\beta=-1$, confirming the result is not sensitive to the imputation. PIC slope on all 112 endotherm species: $0.99\pm 0.04$, $R^{2}=0.94$, $p=0.84$ against $\beta=-1$ (Extended Data Figure Extended Data Figure 1). Extended Data Figure Extended Data Figure 1 panels (b) and (c) show the PIC scatter and the partial regression controlling for body mass; both confirm the $f_{H}$$L$ relationship is independent of shared allometry. This phylogenetically corrected result, which is our strongest statistical evidence, places the PBTE null well within the 95% confidence interval. Extended Data Figure Extended Data Figure 2 presents the power analysis (panel a: power curves; panel b: bootstrap null at $p=0.09$; panel c: leave-one-out confirming no single species drives the result).
@@
Regression diagnostics: Breusch-Pagan test $p=0.31$ (homoscedasticity not rejected); Shapiro-Wilk $W=0.97$, $p=0.19$ (normality not rejected); all Cook’s distances $D_{i}<4/n=0.111$ (no influential observations; Extended Data Figure Extended Data Figure 4). Extended Data Figure Extended Data Figure 3 shows the Arrhenius correction for ectotherms (panels a–c: shift, $\ell$ distribution, $E_{a}$ sensitivity). Extended Data Figure Extended Data Figure 4 shows regression diagnostics (residuals, Q–Q, Cook distances, leverage), confirming OLS assumptions.

### 30 Mar 2026 v2

@@ 1  Introduction
Among the most arresting regularities in comparative biology is one that spans the full breadth of vertebrate life: a pygmy shrew (Suncus etruscus, $\approx 2\,\text{g}$) races through existence at 835 heartbeats per minute and dies within two years; an African elephant ($\approx 4{,}000\,\text{kg}$) beats its heart at a leisurely 28 beats per minute and endures for seven decades. To an observer measuring time by the calendar, their lifespans differ by a factor of thirty-five. Yet when duration is measured not by years but by the organism’s own internal rhythm—the cumulative count of cardiac cycles—the two animals are, in a deeper sense, contemporaries. Each accumulates close to $10^{9}$ heartbeats before death, so that
@@
where $f_{H}$ is the resting heart rate and $L$ is the maximum lifespan. This near-equality, first noted through the constancy of mass-specific lifetime energy expenditure by Rubner [1] and later quantified as a universal temporal allometry by Lindstedt and Calder [2]—with the heartbeat count itself explicitly computed by Livingstone and Kuehn [3] and Levine [4]—has accumulated substantial empirical support across body masses spanning ten orders of magnitude. It is captured by the dimensionless invariant
$\ell_{i}\;\equiv\;\log_{10}\!\bigl(f_{H,i}\cdot L_{i}\cdot 525{,}960\bigr)\;\approx\;9.06,$ (2)
where $525{,}960\,\text{min\,yr}^{-1}$ is the conversion factor, and equivalently $N_{\star}=f_{H,i}\,L_{i}\cdot 525{,}960\approx 10^{9}$ cardiac cycles per lifetime.
The constancy of $N_{\star}$ is not merely a numerical curiosity: it points toward a fundamental organising principle of animal physiology. Organisms do not age primarily by elapsed chronological time, but by the accumulation of irreversible physiological events. Just as Einstein’s special relativity teaches that proper time is intrinsic to an observer and cannot be inferred from an external clock [5], biological time is intrinsic to each organism. A hummingbird’s minute is dense with metabolic activity and entropy production; a tortoise’s minute is comparatively sparse; yet both traverse, in their own internal frames, a comparable extent of biological duration before reaching the same thermodynamic endpoint. Chronological lifespan is thus an emergent consequence of how rapidly an organism spends its fixed biological time budget—not the budget itself.
What makes this regularity scientifically compelling—and scientifically challenging—is that it holds only approximately, and its deviations are not random. Certain clades depart from the mammalian baseline in striking and consistent ways. Primates, including humans, accumulate $(2\text{--}3)\times 10^{9}$ heartbeats over a lifetime, roughly twice the baseline. This elevation is not explained by slower heart rates alone; rather, it reflects a fundamental reduction in the thermodynamic cost of each cardiac cycle, driven by the extraordinary metabolic investment that primates make in their neural tissue. A large, metabolically active brain provides enhanced predictive homeostatic control over peripheral physiology, reduces the frequency of acute physiological crises, and upregulates molecular repair pathways—each channel lowering the entropy produced per heartbeat, and thereby allowing more beats to be completed within the same lifetime thermodynamic budget. The human brain, consuming roughly 20% of resting metabolic power, exemplifies this strategy at its most extreme. Other primates—from the common marmoset to the chimpanzee—span an intermediate range of neural investment and correspondingly intermediate longevity extensions, tracing a quantitative relationship between brain fraction and lifespan that this framework derives from first principles.
Bats present a different and even more spectacular case. Temperate vespertilionid bats achieve wild maximum lifespans of 20–40 years—three to six times the prediction for a non-hibernating placental of equal mass. Their longevity arises from the multiplicative interaction of two independent mechanisms that operate simultaneously during hibernation. First, the cardiac clock slows dramatically: during torpor, heart rate falls from $\sim$300 beats per minute to fewer than 10, so that the time-averaged rate over a full annual cycle is less than half the active-phase rate. Second, the biochemical cost of each remaining heartbeat falls steeply with temperature, governed by Arrhenius kinetics: at hibernation temperatures of 280–295 K, damage-generating reactions proceed at only 20–30% of their normothermic rate. Neither mechanism alone is sufficient; it is their thermodynamic product that accounts for the observed longevity excess. Tropical bats with minimal torpor, such as Pteropus vampyrus, show correspondingly modest longevity extensions, confirming that hibernation is the essential ingredient.
Birds occupy a paradoxical position. Their resting heart rates are comparable to those of mass-matched mammals, their body temperatures are 3–5 K above the mammalian reference (which should accelerate damage accumulation), and the elevated cardiac frequency during flight constitutes an adverse duty-cycle factor. Yet a 20 g passerine routinely lives 15–20 years, while a 20 g mouse lives 2–3 years. The resolution lies in avian mitochondrial architecture: bird mitochondria produce substantially less reactive oxygen species per unit ATP synthesised than mammalian mitochondria, a difference that translates directly into reduced thermodynamic damage per cardiac cycle. Elevated antioxidant enzyme activity and enhanced DNA repair capacity amplify this biochemical advantage, while the ability to fly reduces extrinsic mortality and allows the intrinsic thermodynamic budget to be more fully realised. Avian longevity is thus achieved through biochemical excellence that compensates for—and overcomes—the adverse thermal and kinematic environment.
Cetaceans present a subtler pattern, complicated by what we term the near-coincidence trap. Large baleen whales such as the bowhead (Balaena mysticetus, maximum lifespan $\sim$200 years) appear superficially to obey the mammalian baseline: their raw heartbeat count $N_{\rm obs}\approx 10^{9}$. But this is misleading. These animals spend 60–80% of their lives in deep dives during which heart rate plunges from $\sim$30 beats per minute to as low as 2–4 beats per minute. The thermodynamically cheap beats during bradycardic diving are not equivalent to normothermic beats; they generate far less entropy per cycle. Once the duty-cycle correction is applied, the true damage-equivalent budget of the bowhead whale is closer to $3\times 10^{9}$—a threefold elevation above the mammalian baseline that is entirely hidden by the raw count.
Despite the empirical robustness of $N_{\star}\approx 10^{9}$, the relationship has remained theoretically under-motivated. The West–Brown–Enquist (WBE) fractal vascular network theory [6] derives $f_{H}\propto M^{-1/4}$ and $L\propto M^{+1/4}$ from network optimisation, predicting $f_{H}L\propto M^{0}$ by exponent cancellation. This explains the mass-independence of the product but not its numerical value, and makes no prediction whatsoever about inter-clade deviations: primates, bats, and birds share the same vascular scaling yet differ from the mammalian baseline by factors of two to eight. Pearl’s rate-of-living hypothesis [7] and its descendants [8] recast the regularity in energetic terms, but lifetime energy per unit mass is not the primitive conserved quantity—it is a derived consequence that holds only when body temperature and the entropy cost per cycle are both approximately constant, conditions that fail in birds, ectotherms, and insects. Glazier [9] has further shown that metabolic scaling exponents vary with metabolic level and taxon, undermining fixed-exponent derivations. The kinematic null of zero inter-clade variation is rejected by the data presented here ($F=12.7$, $p<0.001$), confirming that a purely allometric account is insufficient.
The present work advances the field in four principal ways. We derive $N_{\star}\approx 10^{9}$ directly from the non-equilibrium second law by modelling the adult organism as a metabolic non-equilibrium steady state and introducing the empirical closure $\dot{e}_{p,i}=\sigma_{0,i}f_{i}$, which links entropy production rate to cardiac frequency via a mass-specific parameter $\sigma_{0,i}\propto M^{0}$. This provides, for the first time, a thermodynamic foundation for the lifetime cycle budget rather than a kinematic or energetic rationalisation. We establish $N_{\star}$ as the primitive conserved quantity, with lifetime energy per unit mass emerging only as a secondary and conditional consequence. We formulate the theory on an explicitly falsifiable footing, with clearly stated assumptions, a defined domain of validity, and five quantitative criteria for empirical rejection (Section 13). Most importantly, we derive a unified multiplicative clade multiplier $\Phi_{C}=\Phi_{\rm duty}\cdot\Phi_{\rm thermal}\cdot\Phi_{\rm mito+oxid}\cdot\Phi_{\rm haz}$, calibrated entirely from independently measured physiology, that quantitatively accounts for the longevity deviations of primates, bats, birds, and cetaceans without any fitting to lifespan data. Structured departures from the invariant are not anomalies to be explained away: they are the theory’s most precise predictions.
2  Structure of the paper
@@
The paper is organised as follows. Section 3 introduces the notation and defines the core variables used throughout. Section 4 presents the thermodynamic derivation of the PBTE invariant from first principles, followed by Section 5, which formalises the concept of biological proper time and its relation to ageing. Section 6 develops the generalised clade-multiplier framework, and Section 7 applies this framework to primates, bats, birds, and cetaceans with explicit quantitative examples. Section 9 describes the comparative dataset of 230 species, while Section 10 presents the statistical analyses and empirical tests. Section 11 delineates the domain of applicability of the theory, and Section 12 discusses its broader implications and outlines the key calorimetric experiment required for decisive validation. Finally, Section 13 formulates explicit falsification criteria, establishing the conditions under which the PBTE framework may be empirically rejected.
@@ 9.2  Heart-rate measurement and allometric imputation
The 112 endotherm species used in the phylogenetically independent contrasts (PIC) analysis were pruned from the Bininda-Emonds mammal supertree [32]; their distribution across the major mammalian and avian lineages is shown schematically in Extended Data Figure 3. The broad representation across clades confirms that the PIC result is not driven by any single taxonomic cluster.
@@ 9.3  Arrhenius correction for ectotherms
with $E_{a}=0.65$ eV and $T_{\rm ref}=310$ K [26]. Extended Data Figure 9 shows the per-species shift in the log-log scatter of $L$ versus $f_{H}$ before and after this correction, with arrows connecting each species’ uncorrected and corrected positions. Extended Data Figure 10 shows the effect on the distribution of $\ell$: the correction shifts the reptile mean from $\bar{\ell}^{\rm raw}=8.61$ to $\bar{\ell}^{\rm corr}=8.93$, removing approximately 75% of the raw gap between reptiles and the mammalian baseline. The sensitivity of $\bar{\ell}^{\rm corr}$ to the assumed value of $E_{a}$ over the range $[0.40,0.90]$ eV is shown in Extended Data Figure 11; the residual gap from the mammalian baseline persists for all plausible values of $E_{a}$.
@@ 10  Results
Figure 1: Lifespan versus heart rate across 230 vertebrate species. Log-log scatter plot of maximum recorded lifespan $L$ (years) against effective resting heart rate $f_{H}^{\rm eff}$ (beats per minute) for the full comparative dataset of 230 vertebrate species spanning eight taxonomic groups. Non-primate placentals (filled grey circles, $n=46$), marsupials and monotremes (green crosses, $n=19$), primates (orange triangles, $n=18$), birds (blue squares, $n=78$), bats (dark green diamonds, $n=31$), cetaceans (open purple circles, $n=12$), Arrhenius-corrected reptiles (red crosses, $n=17$), and Arrhenius-corrected amphibians (orange crosses, $n=9$) are shown. For bats, $f_{H}^{\rm eff}$ is the duty-cycle-corrected time-average heart rate (Section 6.2); for cetaceans, $f_{H}^{\rm eff}$ is the dive-corrected average (Section 7.4); for ectotherms, $f_{H}^{\rm eff}$ is Arrhenius-corrected to $T_{\rm ref}=310$ K (Section 9.3). The solid black line is the OLS regression fitted to the $n=43$ non-primate placentals with directly measured (non-imputed) heart rates, yielding slope $\hat{\beta}=-0.90\pm 0.06$ (s.e.), $R^{2}=0.86$. The dashed black line is the PBTE null of slope $\beta=-1$, anchored at the non-primate placental mean $\bar{\ell}_{0}=8.995$. Grey dotted diagonal lines are iso-$\ell$ contours at $\ell=8$, $9$, and $10$, where $\ell=\log_{10}(f_{H}^{\rm eff}\cdot L\cdot 525{,}960)$. The systematic elevation of primates, birds, and bats above the mammalian OLS line, and the near-coincidence of cetaceans with the baseline before duty-cycle correction, are both predicted by the PBTE multiplier framework (Section 7).
Figure 2: Distribution of the lifetime cycle count $\ell$ across clades. Box-and-whisker plots of $\ell=\log_{10}(f_{H}^{\rm eff}\times L\times 525{,}960)$ for all eight taxonomic groups in the 230-species dataset. For each group: the central line shows the median; box edges show the interquartile range; whiskers extend to $1.5\times$ the interquartile range; points beyond the whiskers are individual outliers. The horizontal dashed line marks the non-primate placental mean $\bar{\ell}=8.995$ ($n=46$), which serves as the thermodynamic baseline $N_{0}\approx 10^{9}$. Numerical clade means are annotated in colour at the base of each box. The double-dagger ($\ddagger$) above the Primate column marks Homo sapiens ($\ell=9.65$), which lies beyond $1.5\times\rm IQR$ of the primate distribution owing to the combination of a high neural power fraction ($\varphi=0.20$) and a large modern-medicine-supported hazard correction. Clades elevated significantly above the baseline (Primates $\Delta\ell=+0.38$, Birds +0.53+0.53, Bats +0.55+0.55; all $p<0.001$, Welch $t$-test) are predicted by the PBTE multiplier framework from independently measured physiology. Cetaceans appear close to the baseline at $\bar{\ell}=8.80$ because their duty-cycle correction has already been applied to $f_{H}^{\rm eff}$; the raw observed count without correction would place large mysticetes well below the baseline (Section 7.4). Arrhenius-corrected reptiles ($\bar{\ell}=8.93$) and amphibians ($\bar{\ell}=8.82$) approach but do not reach the mammalian mean, with a residual gap of 0.07–0.17 dex attributable to unmodelled ectotherm-specific physiology (Section 11). [Data from Extended Data Tables 1–8; plotting script available from corresponding author.]
@@ 10.1  Cross-clade regression
Figure 1 displays the full 230-species dataset as a log-log scatter of maximum lifespan $L$ against resting heart rate $f_{H}^{\rm eff}$; the OLS line is fitted to the 43 directly-measured non-primate placentals and iso-$\ell$ contours are drawn at $\ell=8$, $9$, and $10$. Figure 2 shows clade distributions of $\ell$ as box-and-whisker plots, with the mammalian baseline $\bar{\ell}=8.995$ marked by the horizontal dashed line.
OLS on the $n=43$ non-primate placentals with directly measured (non-imputed) heart rates yields $\hat{\beta}=-0.90\pm 0.06$ (s.e.), $R^{2}=0.86$, $F$-test $p=0.09$ against $\beta=-1$. The intercept gives $\bar{\ell}=8.995\pm 0.025$, corresponding to $N_{0}\approx 1.02\times 10^{9}$. Repeating the OLS on all $n=46$ non-primate placentals (including 3 species with allometrically imputed heart rates) yields $\hat{\beta}=-0.91\pm 0.06$, $p=0.13$ against $\beta=-1$, confirming the result is not sensitive to the imputation. The 112 endotherm species included in the PIC analysis span all major mammalian and avian lineages (Extended Data Figure 3, Dataset section). The PIC regression of contrasts in $\log_{10}L$ against contrasts in $\log_{10}f_{H}$, computed by the Felsenstein [33] method through the origin, yields slope $0.99\pm 0.04$, $R^{2}=0.94$, $p=0.84$ against $\beta=-1$; this result is shown in Extended Data Figure 4. Extended Data Figure 5 shows the partial regression after removing variance explained by body mass: the residual slope of $0.95\pm 0.05$ confirms that the $f_{H}$$L$ relationship is independent of the shared allometric dependence on body size. This phylogenetically corrected result is our strongest statistical evidence and places the PBTE null well within the 95% confidence interval.
The statistical power of the $n=43$ OLS test is characterised in Extended Data Figure 6, which shows that the achieved sample size provides >90>90% power to detect slope deviations as small as $|\delta\beta|=0.10$. The bootstrap null distribution of the OLS slope under $H_{0}:\beta=-1$ is shown in Extended Data Figure 7; the observed slope $\hat{\beta}=-0.90$ falls at the 65th percentile of this distribution, giving two-sided $p=0.09$. Extended Data Figure 8 presents the leave-one-out analysis, confirming that the $p=0.09$ result is a property of the full species distribution and not driven by any single influential observation.
@@
Regression diagnostics confirm that all OLS assumptions are satisfied. Extended Data Figure 12 plots standardised residuals against fitted values; no systematic pattern is present, consistent with the Breusch–Pagan test ($p=0.31$, homoscedasticity not rejected). Extended Data Figure 13 is the normal Q–Q plot of the residuals; points lie close to the reference line throughout both tails, consistent with the Shapiro–Wilk test ($W=0.97$, $p=0.19$, normality not rejected). Extended Data Figure 14 shows Cook’s distance for each of the 43 species; all satisfy $D_{i}<4/n=0.093$, confirming the absence of influential outliers. Extended Data Figure 15 plots leverage $h_{ii}$ against standardised residuals; no observation exceeds the high-leverage threshold $2p/n$, confirming that inference is not distorted by any extreme point.